OpenStudy (anonymous):

Linear Algebra. Let A and B be 5x5 matrices, such that B is invertible. Suppose that AB =-BA.Show that A is not invertible.

OpenStudy (zarkon):

look at the determinant

OpenStudy (anonymous):

\[ A B = - B A\\ Det( AB) = (-1)^5 Det(B A)\\ Det(A) Det(B) = -Det(B)Det(A)\\ Det(A) = - Det(A)\\ \] Since \[ Det(B) \ne 0 \] Can you conclude now that A is not invertible?

OpenStudy (anonymous):

The above is true for 5 replaced by any odd integer.